OpenStudy (anonymous):

Find the derivative at each critical point and determine the local extreme values of y=x^2/3(x^2-4); x≥0.

OpenStudy (tkhunny):

Okay, you should get working on that derivative. What's your plan?

OpenStudy (anonymous):

Not a clue honestly. I tried to derive it using the product rule and it's safe to safe that I failed at that. So now I figured to distribute the x^2/3 so working on that now!

OpenStudy (tkhunny):

You're saying the Quotient Rule didn't cross your mind? It is a division problem.

OpenStudy (tkhunny):

Wait, is it \(\dfrac{x^{2}}{3}(x^{2} - 4)\) or maybe \(\dfrac{x^{2}}{3(x^{2} - 4)}\)?

OpenStudy (tkhunny):

It may help just to get the 3 out of the way. \(\dfrac{d}{dx}\dfrac{x^{2}}{3}(x^{2} - 4) = \dfrac{1}{3}\dfrac{d}{dx} x^{2}(x^{2} - 4)\) You can work on the parts with variables and just keep the 1/3 along for the ride.

OpenStudy (anonymous):

I see it as a multiplication problem. I'm not quite sure how you got the 3 in the denominator.

OpenStudy (tkhunny):

Isn't that what "/3" means?

OpenStudy (anonymous):

Yes that is true, I don't know how to write it out all fancy like you. /3 is part of 2/3. x raised to the two thirds. It's a fractional exponent that just belongs to the x.

OpenStudy (anonymous):

(x^2/3)(x^2-2) maybe makes more sense?

OpenStudy (anonymous):

(x^2/3)(x^2-4) is actually the correct equation.

OpenStudy (tkhunny):

Please remember your Order of Operations. x^2/3 = \(x^{2}/3 = \dfrac{x^{2}}{3}\) x^(2/3) = \(x^{2/3}\) Multiplying through should work just fine. [x^(2/3)]*(x^2 - 2) = x^(8/3) - 2x^(2/3) Okay, now what?